Convert 30 years to a) seconds and b) day


1
Expert's answer
2021-03-31T11:22:35-0400

(a)

"30\ y\cdot\dfrac{365\ d}{1\ y}\cdot\dfrac{24\ h}{1\ d}\cdot\dfrac{3600\ s}{1\ h}=9.461\cdot10^8\ s."

(b)

"30\ y\cdot\dfrac{365\ d}{1\ y}=10950\ d."